# FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering

### v3

@@ Abstract
Federated learning (FL) has become a key component in various language modeling applications such as machine translation, next-word prediction, and medical record analysis. These applications are trained on datasets from many FL participants that often include privacy-sensitive data, such as healthcare records, phone/credit card numbers, login credentials, etc. Although FL enables computation without necessitating clients to share their raw data, determining the extent of privacy leakage (PL) in federated language models is not straightforward. Moreover, existing attacks aim to extract data regardless of how sensitive it is. To fill this research gap, we introduce two novel findings in leaking privacy-sensitive user data from federated large language models. Firstly, we make a key observation that model snapshots from the intermediate rounds in FL can cause greater privacy leakage than the final trained model. Secondly, we identify that PL can be aggravated by tampering with a model’s selective weights that are specifically responsible for memorizing the sensitive training data. We show how a malicious client in FL can leak the privacy-sensitive data of some other clients in FL even without any cooperation from the server. Our best-performing method improves the membership inference recall by 29% and achieves up to 71% private data reconstruction, evidently outperforming existing attacks with stronger assumptions of adversary capabilities.
1 Introduction
@@
Large language models (LLMs) are playing a pivotal role in various fundamental natural language processing (NLP) tasks [30, 32]. Recently, these LLMs have been trained on extremely large text corpora [21] and grown to billions of parameters [6, 68]. These pre-trained LLMs are then adapted to various downstream tasks by fine-tuning on domain-specific datasets, often on privacy-sensitive datasets such as in-house emails [26], code repositories [^1] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset, clinical health data [28], etc. Given the privacy-sensitive nature of these datasets, collecting such data by a centralized entity to fine-tune an LLM poses an unprecedented level of privacy concern. To solve such issues, Federated learning (FL) [61, 64, 67] has emerged as a widely-used technology. FL allows language models to be trained by exchanging only model updates, without the need for a central entity to collect all the training data. However, recent studies [71, 80] have shown that carefully crafted attacks can invert the model updates sent by FL clients and recover private information. To make matters worse, large LLMs tend to output complete sequences verbatim from their training data [9, 7], which by itself is a major privacy concern. Given the increasing use of FL technologies in language models, it is now of paramount importance to investigate the implications of privacy leakage (PL) in such settings.
@@
Most prior work investigating the privacy of language models focus on centralized settings [9, 8, 23] and/or rely on external adversaries [24, 76]. However, the knowledge of privacy implications in centralized learning scenarios often does not transfer to federated learning due to the fundamental differences in their training protocols. For instance, in a centralized setting, the training process sees all the data points in every epoch. In contrast, federated learning uses asynchronous aggregation [52] where the server selects a subset of clients with a fraction of data in each round of training. The only few works [19, 24, 3] that focus on privacy attacks in federated language models aim to extract data indiscriminately as opposed to devising attacks to extract specific types of privacy-sensitive data (e.g., health-related data) of the FL clients/users. To fill this research gap, in this paper, we investigate the leakage of users’ private data in FL settings and design novel privacy attacks that maximally exploit the adversary’s capability to leak sensitive data of FL clients, particularly the types of data the adversary is interested in. For instance, email service providers (e.g., Gmail, Outlook) may train an FL model for their auto-complete feature that learns from the users’ emails without requiring direct access to email data, thus providing the privacy promised by FL. However, users may share various types of privacy-sensitive data such as SSNs or credit card numbers in their emails. It would be a serious privacy concern if one user of such a federated setting could extract sensitive information about other users. Our attacks show that this is indeed possible where the attacker targets to extract some sensitive data from other users. Such extracted data could then be used to undertake serious security and privacy crimes, e.g., abusing extracted credit card information, performing SSN reverse lookup, and then impersonating the victim to banks/other institutions.
@@
Understanding the extent of PL in FL-based language models is complex and challenging. This is because previous research on forgetting analysis [29] asserts that deep language models tend to forget the early training samples gradually if those are not repeated much afterward. Additionally, as seen in studies on catastrophic forgetting [20, 31] during iterative training, models could lose information about earlier samples as they see new training samples. Since only a subset of the clients participate in each round of FL, the above observations are inherently true for FL settings. The federated averaging (FedAvg) algorithm [45] also reduces the memorization of such data points as the training heads toward convergence [66]. These observations imply that the language model tends to forget any client’s local data as training proceeds and tries to generalize more on the overall distribution of the entire training corpus. This could potentially make it difficult for an attacker to extract a client’s private data from the final model in FL. However, existing research has not yet explored the exploitation of intermediate model snapshots for privacy attacks in federated language models. Hence, to fill this research gap, we exploit the model snapshots from intermediate rounds of FL in our attack design, where the participating clients have privacy-sensitive text data, e.g., phone numbers, SSNs, passwords, etc. in their local datasets. We show how a malicious FL participant can cause leakage of some other FL participants’ sensitive data with and even without the server’s cooperation by maximizing the memorization of those intermediate model snapshots.
@@
Besides exploiting this intrinsic vulnerability of the FL configuration in our attack design, we further investigate how the attacker could enhance the model’s memorization of privacy-sensitive training data. The proposed methods stem from our crucial observation that some weights of an LLM are more susceptible to sensitive data points than others due to their out-of-distribution nature. Specifically, the model heads and nearby layers in the transformer blocks have the strongest influence [49]. Therefore, strategically modifying these selective weights of an LLM can effectively increase its memorization of privacy-sensitive data. Based on this observation, we designed two methods for tuning the model weights and piloting them toward stronger memorization. The first one is called selective weights optimization (SWO), which takes into account the difference of selective weights between a fine-tuned (with privacy-sensitive data) model and another reference model, and then maximizes this difference. The second approach, named weights transformation learning (WTL), utilizes an end-to-end supervised algorithm to learn the hidden pattern of how those selective weights mutate from a reference model to a fine-tuned model. The attacker may apply these techniques either actively during fine-tuning (referred to as dynamic attack) or passively after the fine-tuning is done (referred to as static attack). Our dynamic mode of the attack closely resembles the recent line of works using model poisoning for PL [78, 4, 18, 72, 17]. However, most of these existing works consider only centralized learning scenarios and poison the pre-trained model beforehand [72, 17, 41]. They also presume strong adversarial capabilities, e.g., assuming the server to be an active adversary [51] or the adversary changing the model’s architecture [18]. In contrast, our static attack, with no model poisoning or server cooperation, results in up to 63% sensitive data reconstruction, while our dynamic attack, involving active weight tampering during FL fine-tuning, improves the reconstruction rate up to 71%. Both attack modes are highly effective on open-source LLMs, including Gemma [65], Llama-2 [68], GPT-2 [59], and BERT [13]. We also emphasize that our attacks are not domain-specific. They are applicable as long as the sensitive texts are somewhat out-of-distribution in nature, such as any domain-specific jargon, medical prescriptions, and forensic test results.
Although previous works on PL attacks exploited the outlier nature of private texts [8, 26, 42], ours is the first work to maximize data memorization of a federated language model by tampering with its selective weights to increase the privacy leakage. Our main contributions are as follows:
@@
❒ We propose a novel attack technique that exploits model snapshots from intermediate rounds in FL to leak the privacy-sensitive text data of the FL users.
@@
❒ We identify the weights largely responsible for memorizing privacy-sensitive training data and introduce two novel methods- SWO and WTL to maximize the model’s memorization by tampering with those selective weights.
@@
❒ Without any strong assumption on the adversary’s capability, our proposed strategies perform both membership inference and reconstruction attacks on clients’ private data and significantly outperform existing attacks [19, 24].
❒ Our attack methods are highly effective on auto-regressive and masked language modeling tasks with Gemma, Llama-2, GPT-2, and BERT LLM, using both artificial canaries and real-life privacy-sensitive data.
2 Related Work
2.1 Privacy Leakage Attacks in LLMs
@@
The privacy risks of LLMs in centralized and federated learning settings have been extensively studied in prior works. The seminal work by Carlini et al. [9] proposes an attack to retrieve public internet texts by querying a pre-trained GPT-2 model. Our attack, however, aims to extract privacy-sensitive data when fine-tuning an LLM in federated learning. In the domain of RNN and LSTM models, Carlini et al. [8] highlight the problem of unintended memorization. They insert canaries into the training dataset and measure their leakage probability with a metric called ‘exposure’. We use this exposure metric in our work to show the effectiveness of our novel attack strategies. Mireshghallah et al. [49] study the effect of fine-tuning on memorization and show that fine-tuning the parameters higher in the LLM architecture results in the most data leakage. Our experiments in federated learning align with this observation, and we manipulate these selective weights of an LLM that are most susceptible to private data for producing greater memorization. Kim et al. [33] and Lukas et al. [42] have proposed algorithms to evaluate to what extent information memorized by the LM constitutes sensitive personally identifiable information (PII) and whether existing defenses are sufficient to prevent leakage. Although our objectives in this work are similar, all of the aforementioned works have been done considering a non-FL threat model, whereas our attack methods ensure maximum user data leakage by exploiting the model parameters in FL settings.
2.2 Privacy Leakage By Poisoning
@@
Recent research has explored model and data poisoning to leak privacy, though most studies do not focus on language modeling [78, 4]. Wen et al. [72], Feng et al. [17], and Liu et al. [41] manipulate the weights of pre-trained LLMs to cause privacy leakage (PL) during fine-tuning. However, these attacks do not apply to federated learning (FL). On the other hand, Fowl et al. [18] consider the server in FL as an adversary that actively poisons model architecture and parameters sent to users. Such attacks are easily detectable in pre-trained LLMs, where the architecture is well-known to clients. In contrast, our static attack method can passively leak the privacy of FL users after fine-tuning without poisoning the pre-trained model, thus fully preserving the utility of the model. Additionally, our dynamic attack can further enhance PL by actively tampering with the model’s weights during FL fine-tuning, where the attacker is one of the FL clients, not the server.
2.3 Privacy Leakage in Federated Learning
@@
Decepticons [19] and FILM [24] are attacks in federated learning settings. Decepticons consider a setting where an active adversary server maliciously alters the model state before sending them to the clients. FILM considers an external adversary that eavesdrops on the communication channel between the server and clients to reconstruct sentences from gradients. Our attack differs from them as we neither consider the server as an active adversary nor assume the presence of a secure communication channel. Section 5.6 presents a detailed comparison of our attack with Decepticons and FILM. The LAMP [3] attack considers an attacker that observes the gradients in the FedSGD algorithm and inverts them into natural text using some discrete optimization algorithms. Nasr et al. [50] also did something similar where their method stands on the assumption that gradient updates are available to the users and server, like in the case of FedSGD. In contrast, we consider FedAvg where the gradient information is not shared across clients/servers. Privacy breaches through gradient inversion are quite trivial at present for both language models and other deep learning models [27]. In our paper, we are dealing with a more challenging threat model where gradient inversion/ manipulation attacks are not possible. In addition, all these attacks on federated language models attempt to extract data indiscriminately with no extra focus on privacy-sensitive data. Hence, we come up with novel attack strategies to breach private user data in FL.
@@
3 Impact of Privacy-Sensitive Data on Fine-tuning Large Language Models
@@
To carry out different downstream applications, large language models (LLMs) such as GPT-2 [59] and BERT [13] undergo a fine-tuning step that modifies their pre-trained weights according to those specific tasks [11]. However, privacy-sensitive texts have specific patterns and often contain out-of-distribution tokens, e.g., digits (in phone or credit card numbers), special characters, and out-of-vocabulary words (login credentials or physical addresses). To understand the impact of such privacy-sensitive data on fine-tuning, we investigate whether the changes in the pre-trained weights after fine-tuning the LLMs with such sensitive data show any differentiating patterns when compared to the weight changes after fine-tuning with regular English texts. Our key observation from this investigation is that there indeed exists a remarkable difference between the patterns in these two cases. Figure 1(a) depicts the norms of weight changes in all 12 transformer blocks, i.e., multi-layer perception (MLP) and self-attention layers after fine-tuning a GPT-2 (base) model $G$ with regular English sentences $\mathcal{D}_{reg}$. Here, we observe a uniform impact on the weights across different layers, $L$, which suggests that the model’s adaptation after fine-tuning is evenly distributed throughout the architecture. However, a distinct pattern emerges when fine-tuning $G$ with privacy-sensitive texts $\mathcal{D}_{sen}$ such as phone numbers, email addresses, login credentials, credit card numbers, medical prescriptions, etc. Only a subset of layers, specifically the model heads and nearby layers in the transformer blocks, denoted as $L_{\star}\subset L$, undergo a substantial transformation, while the remaining layers, denoted as $L_{\ominus}=L\setminus L_{\star}$, exhibit minimal/regular changes. Figure 1(b) shows the transformation of the MLP and self-attention layers of GPT-2 after fine-tuning with $D_{sen}$ (Refer to Figure 8 for Gemma). We experiment with four different language models - Gemma, Llama-2, GPT-2, and BERT - using various types of sensitive data and find that as long as these sensitive texts have an out-of-distribution characteristic with respect to the training data, each of these LMs shows such disproportionate changes in their internal weights. This crucial observation supports the following hypothesis:
@@
Here, $W(f_{G}(D))$ denotes the weights of model $G$ fine-tuned with dataset $D$. We observe that $\|\Delta W_{\star}\|>>\|\Delta W_{\ominus}\|$. In contrast, when fine-tuning with regular texts, we observe $\|\Delta W_{\star}\|\sim\|\Delta W_{\ominus}\|$. It is pertinent to mention that the embedding layer of an LM is solely responsible for converting input tokens into vectors with positional information, while the other components, including layer normalization, do not actively impact fine-tuning [59]. Hence, we only concentrate on MLP and self-attention layers of transformer blocks.
@@
This observation indicates that weights of certain layers ($L_{\star}$ marked with green rectangles in Figures 1(a) and 1(b)) within an LLM are more susceptible to capturing and memorizing privacy-sensitive information. We also scrutinize the direction/pattern of changes in those selective weights caused by out-of-distribution samples in the fine-tuning dataset. Then we ask the most important question- if we further increase the amplitude of those particular weight changes along the same direction, does it make the model more biased towards the out-of-distribution data? Later, in Section 4.2.2 where we discuss our Maximizing Data Memorization (MDM) techniques, we illustrate that the answer to this question is ‘Yes’, and we can strategically focus on modifying the weights of these sensitive layers to enhance the model’s memorization of private data. But that is not all. With this also comes the salient question- If we amplify the weight changes in such a manner, does that affect the model utility? We also convincingly answer this question in Section 4.2.2 by showing that such modification hardly impacts the model’s learning of regular English texts.
@@
We qualitatively define memorization $\mathbf{M}$ as a property of the language model in exposing particular training data points during inference. Quantitative measure of memorization $\mathbf{M}$ on $\mathcal{D_{\text{sen}}}$ includes but are not limited to data reconstruction [8], membership inference [44], and exposure, which we discuss in detail in Section 4.5.
@@
Table 1: A high-level taxonomy of our attack methods
4 Attack Methodology
4.1 Threat model

### v4

# Gradient-Free Privacy Leakage in Federated Language Models through Selective Weight Tampering

@@ Abstract
Federated learning (FL) has become a key component in various language modeling applications such as machine translation, next-word prediction, and medical record analysis. These applications are trained on datasets from many FL participants that often include privacy-sensitive data, such as healthcare records, phone/credit card numbers, login credentials, etc. Although FL enables computation without necessitating clients to share their raw data, existing works show that privacy leakage is still probable in federated language models. In this paper, we present two novel findings on the leakage of privacy-sensitive user data from federated large language models without requiring access to gradients. Firstly, we make a key observation that model snapshots from the intermediate rounds in FL can cause greater privacy leakage than the final trained model. Secondly, we identify that a malicious FL participant can aggravate the leakage by tampering with the model’s selective weights that are responsible for memorizing the sensitive training data of some other clients, even without any cooperation from the server. Our best-performing method [^1] Our source code: https://figshare.com/s/8636bd193495193c0fe4 increases the membership inference recall by 29% and achieves up to 71% private data reconstruction, evidently outperforming existing attacks that consider much stronger adversary capabilities. Lastly, we recommend a balanced suite of techniques for an FL client to defend against such privacy risk.
I Introduction
@@
Large language models (LLMs) are playing a pivotal role in various fundamental natural language processing (NLP) tasks [joshi2017triviaqa, kieuvongngam2020automatic]. Recently, these LLMs have been trained on extremely large text corpora [gao2021pile] and grown to billions of parameters [brown2020language, touvron2023llama]. These pre-trained LLMs are then adapted to various downstream tasks by fine-tuning on domain-specific datasets, often on privacy-sensitive datasets such as in-house emails [huang2022large], code repositories [bigcode], clinical health data [jagannatha2021membership], etc. Given the privacy-sensitive nature of these datasets, collecting them by a centralized entity to fine-tune an LLM raises significant privacy concerns. To solve such issues, federated learning [shokri2015privacy, sui2020feded, tian2022fedbert] has emerged as a widely-used technology. FL allows language models to be trained by exchanging only model updates, without requiring a central entity to collect all the training data. However, recent studies [wang2019beyond, zhao2020idlg] have shown that carefully crafted attacks can invert the model updates sent by FL clients and recover private information. To make matters worse, LLMs tend to output complete sequences verbatim from their training data [carlini2021extracting, carlini2023quantifying], which by itself is a major privacy concern. Given the increasing use of FL technologies in language models, it is now of paramount importance to investigate the implications of privacy leakage in such settings.
@@
Most prior work investigating the privacy of language models focuses on centralized settings [carlini2021extracting, carlini2019secret, gu2023towards] and/or relies on external adversaries [gupta2022recovering, yuan_etal]. However, the knowledge of privacy implications in centralized learning scenarios often does not transfer to federated learning due to the fundamental differences in their training protocols. For instance, in a centralized setting, the training process sees all the data points in every epoch. In contrast, federated learning uses asynchronous aggregation [nishio2019client], where the server selects a subset of clients, each with a fraction of the data, in each round of training. There are a few works that focus on privacy attacks in federated language models, but they aim to extract data from FL users indiscriminately [fowl2022decepticons, gupta2022recovering, lamp] rather than devising more impactful attacks that specialize in extracting privacy-sensitive data, e.g., PII (Personally Identifiable Information). To fill this research gap, in this paper, we investigate the leakage propensity of users’ privacy-sensitive data that an adversary is particularly interested in. For instance, email service providers (e.g., Gmail, Outlook) may train an FL model for their auto-complete feature that learns from the users’ emails without requiring direct access to email data, thus providing the privacy promised by FL. However, users may share various types of privacy-sensitive data, such as SSNs or credit card numbers, in their emails. It would be a serious privacy concern if one user of such a federated setting could extract sensitive information about other users. Our findings show that this is indeed possible, where the attacker is curious about some specific type of sensitive data to extract from other users. Such extracted data could then be used to undertake serious security and privacy crimes, e.g., abusing extracted credit card information, performing SSN reverse lookup, and then impersonating the victim to banks/other institutions.
@@
Learning algorithms like FedSGD [mcmahan2016communication], which allow unrestricted access to gradients, make the privacy-breach challenge quite trivial for the adversary, as many gradient-inversion attacks [leakage_gradients, huang2021evaluating, jiang2022comprehensive] have become the state-of-the-art in red-teaming against an ML model. However, in this work, we deal with a much tighter FL implementation, i.e., federated averaging (FedAvg), where the gradient information is kept private and only weights are shared [mcmahan2016communication]. Previous research on information leakage [hatamizadeh2023gradient] and inversion attacks [huang2021evaluating, elkordy2022much] in FL concludes that FedAvg is significantly harder to attack or leaks less privacy than FedSGD. Nevertheless, understanding the extent of privacy leakage in FL-based language models is complex and challenging. This is because previous research on forgetting analysis [jagielski2022measuring] shows that deep language models tend to gradually forget early training samples if they are not repeated much afterward. Additionally, as seen in studies on catastrophic forgetting [french1999catastrophic, kemker2018measuring], during iterative training, models can lose information about earlier samples as they encounter new ones. Since only a subset of the clients participates in each round of FL, the above observations are inherently true for FL settings. The FedAvg algorithm also reduces the memorization of such data points as the training heads toward convergence [thakkar2020understanding]. These observations imply that the model tends to forget any client’s local data as training proceeds and tries to generalize more on the overall distribution of the entire training corpus. This could potentially make it difficult for an attacker to extract a client’s private data from the final model in FL. However, existing research has not yet explored the privacy vulnerabilities related to the intermediate model snapshots in federated language models. Hence, in our attack design, we exploit the model snapshots from intermediate rounds of FL, where the participating clients have privacy-sensitive text, e.g., phone numbers, SSNs, passwords, etc. in their local datasets. We show how a malicious FL participant can retrieve some other FL participants’ sensitive data even without the server’s cooperation by maximizing the memorization of those intermediate model snapshots.
@@
Besides exploiting the FL configuration’s intrinsic vulnerability in our attack design, we further investigate how an attacker could enhance the model’s memorization of privacy-sensitive training data. The proposed methods stem from the crucial observation that some weights of an LLM are more susceptible to sensitive data points than others due to their out-of-distribution nature. Specifically, the MLP sub-layers in the last few transformer blocks have the strongest influence [mireshghallah2022memorization, chen2024learnable]. Therefore, strategically adjusting these selective weights in an LLM can effectively improve its ability to memorize privacy-sensitive data. Based on this observation, we designed two methods for tuning the model weights and piloting them toward stronger memorization. The first one is a zeroth-order method called selective weights optimization (SWO), which takes into account the difference of selective weights between a fine-tuned (with privacy-sensitive data) model and another reference model, and then maximizes this difference. The second approach, named weights transformation learning (WTL), utilizes an end-to-end supervised algorithm to learn the hidden pattern of how those selective weights mutate from a reference model to a fine-tuned model. The attacker may apply these techniques either actively during fine-tuning (referred to as dynamic attack) or passively after the fine-tuning is done (referred to as static attack). Our dynamic mode of the attack closely resembles the recent line of works using model poisoning for Privacy Leakage [zhang2023agrevader, 10190537, fowl2021robbing, wen2024privacy, feng2024privacy]. However, most of these existing works consider only centralized learning scenarios and poison the pre-trained model beforehand [wen2024privacy, feng2024privacy, liu2024precurious]. They also presume strong adversarial capabilities, e.g., assuming the server to be an active adversary [nguyen2023active], accessing the gradients [10190537], or changing the model’s architecture [fowl2021robbing]. In contrast, our static attack, with no model poisoning or server cooperation, results in up to 63% sensitive data reconstruction, while our dynamic attack, involving active weight tampering during FL fine-tuning, improves the reconstruction rate up to 71%. Both attack modes are highly effective on open-source LLMs, including Gemma [team2024gemma], Llama-2 [touvron2023llama], GPT-2 [radford2019language], and BERT [devlin-etal-2019-bert]. We also emphasize that our attacks are not domain-specific. They are applicable as long as the sensitive texts are somewhat out-of-distribution in nature, such as any domain-specific jargon, medical prescriptions, and forensic test results. Although previous works on Privacy Leakage attacks exploited the outlier nature of private texts [carlini2019secret, huang2022large, lukas2023analyzing], ours is the first work to maximize data memorization of a federated language model by tampering with its selective weights to increase the privacy leakage. Finally, we experiment with possible privacy defenses, including differential-privacy [dwork_roth_dp] , regularization [wu2023depn], and scrubbing [chowdhury2021adversarial], and recommend the best possible combination of methods for an FL client to defend against such privacy leakage. Our main contributions are as follows:
@@
❒We propose a gradient-free novel attack framework that exploits model snapshots from intermediate rounds in FL to leak the privacy-sensitive text data of the FL users.
@@
❒We introduce two novel methods- SWO and WTL to maximize LLM’s memorization by tampering with selective weights, which are largely responsible for memorizing privacy-sensitive data.
@@
❒Our proposed framework functions as a pre-attack routine, bolstering both membership inference and reconstruction attacks on clients’ private data, significantly outperforming existing attacks.
❒We empirically analyze possible defenses and recommend the best suite of methods to mitigate such privacy threats.
II Related Work
II-A Privacy Leakage Attacks in LLMs
@@
The privacy risks of LLMs in centralized and federated learning settings have been extensively studied in prior works. The seminal work by Carlini et al. [carlini2021extracting] proposes an attack to retrieve public internet texts by querying a pre-trained GPT-2 model. Our attack, however, aims to extract privacy-sensitive data when fine-tuning an LLM in federated learning. In the domain of RNN and LSTM models, Carlini et al. [carlini2019secret] highlight the problem of unintended memorization. They insert canaries into the training dataset and measure their leakage probability with a metric called ‘exposure’. We use this exposure metric in our work to show the effectiveness of our novel attack strategies. Mireshghallah et al. [mireshghallah2022memorization] study the effect of fine-tuning on memorization and show that fine-tuning the parameters higher in the LLM architecture results in the most data leakage. Our experiments in federated learning align with this observation, and we manipulate these selective weights of an LLM that are most susceptible to private data for producing greater memorization. Kim et al. [kim2023propile] and Lukas et al. [lukas2023analyzing] have proposed algorithms to evaluate to what extent information memorized by the LM constitutes sensitive personally identifiable information (PII) and whether existing defenses are sufficient to prevent leakage. Although our objectives in this work are similar, all of the aforementioned works have been done considering a non-FL threat model, whereas our attack methods ensure maximum user data leakage by exploiting the model parameters in FL settings.
II-B Privacy Leakage By Poisoning
@@
Recent research has explored model and data poisoning to leak privacy, though most studies do not focus on language modeling [zhang2023agrevader, 10190537]. Wen et al. [wen2024privacy], Feng et al. [feng2024privacy], and Liu et al. [liu2024precurious] manipulate the weights of pre-trained LLMs to cause privacy leakage during fine-tuning. However, these attacks do not apply to FL. On the other hand, Fowl et al. [fowl2021robbing] consider the server in FL as an adversary that actively poisons model architecture and parameters sent to users. Such attacks are easily detectable in pre-trained LLMs, where the architecture is well-known to clients. In contrast, our static attack method can passively leak the privacy of FL users after fine-tuning without poisoning the pre-trained model, thus fully preserving the utility of the model. Additionally, our dynamic attack can further enhance Privacy Leakage by actively tampering with the model’s weights during FL fine-tuning, where the attacker is one of the FL clients, not the server.
II-C Privacy Leakage in Federated Learning
@@
Decepticons [fowl2022decepticons] and FILM [gupta2022recovering] are attacks in federated learning settings. Decepticons consider a setting where an active adversary server maliciously alters the model state before sending them to the clients. FILM considers an external adversary that eavesdrops on the communication channel between the server and clients to reconstruct sentences from gradients. Our attack differs from theirs as we neither consider the server as an active adversary nor assume their access to gradients. Section VI-C presents a detailed comparison of our attack with Decepticons and FILM. The LAMP [lamp] attack considers an attacker that observes the gradients in the FedSGD algorithm and inverts them into natural text using some discrete optimization algorithms. Nasr et al. [nasr2019comprehensive] also did something similar where their method stands on the assumption that gradient updates are available to the users and server, like in the case of FedSGD. In contrast, we consider FedAvg where the gradient information is not shared across clients/servers. Privacy breaches through gradient inversion are quite trivial at present for both language models and other deep learning models [huang2021evaluating]. In our paper, we are dealing with a more challenging threat model where gradient inversion/ manipulation attacks are not possible. In addition, all these attacks on federated language models try to extract data indiscriminately with no extra focus on privacy-sensitive data. Hence, we come up with novel attack strategies to breach private user data in FL.
@@
III Impact of Privacy-Sensitive Data on Fine-tuning Large Language Models
@@
To carry out different downstream applications, LLMs such as GPT-2 [radford2019language] and BERT [devlin-etal-2019-bert] undergo a fine-tuning step that modifies their pre-trained weights according to those specific tasks [chung2022scaling]. However, privacy-sensitive texts have specific patterns and often contain out-of-distribution tokens, e.g., digits (in phone or credit card numbers), special characters, and out-of-vocabulary words (login credentials or physical addresses). To understand the impact of such privacy-sensitive data on fine-tuning, we investigate whether the changes in the pre-trained weights after fine-tuning the LLMs with such sensitive data show any differentiating patterns when compared to the weight changes after fine-tuning with regular English texts. Our key observation from this investigation is that there indeed exists a remarkable difference between the patterns in these two cases. Figure 1(a) depicts the norms of weight changes in all 12 transformer blocks, i.e., multi-layer perceptron (MLP) and self-attention layers after fine-tuning a GPT-2 (base) model $G$ with regular English sentences $\mathcal{D}_{reg}$. Here, we observe a uniform impact on the weights across different layers, $L$, which suggests that the model’s adaptation after fine-tuning is evenly distributed throughout the architecture. However, a distinct pattern emerges when fine-tuning $G$ with privacy-sensitive texts $\mathcal{D}_{sen}$ such as phone numbers, email addresses, login credentials, credit card numbers, medical prescriptions, etc. Only a subset of weights, specifically the MLP sub-layers in the last few transformer blocks, denoted as $L_{\star}\subset L$, undergo a substantial transformation, while the remaining layers, denoted as $L_{\ominus}=L\setminus L_{\star}$, exhibit minimal/regular changes. This exact observation is also supported by Chen et al. [chen2024learnable]. Figure 1(b) shows the transformation of the MLP and self-attention layers of GPT-2 after fine-tuning with $D_{sen}$ (Refer to Figure 7 for Gemma). We experiment with four different language models - Gemma, Llama-2, GPT-2, and BERT - using various types of sensitive data and find that as long as these sensitive texts have an out-of-distribution characteristic with respect to the training data, each of these LMs shows such disproportionate changes in their internal weights. This crucial observation supports the following hypothesis:
@@
Here, $W(f_{G}(D))$ denotes the weights of model $G$ fine-tuned with dataset $D$. We observe that $\|\Delta W_{\star}\|>>\|\Delta W_{\ominus}\|$. In contrast, when fine-tuning with regular texts, we observe $\|\Delta W_{\star}\|\sim\|\Delta W_{\ominus}\|$.
@@
This observation indicates that weights of certain layers ($L_{\star}$ marked with green rectangles in Figures 1(a) and 1(b)) within an LLM are more susceptible to capturing and memorizing privacy-sensitive information. We also scrutinize the direction/pattern of changes in those selective weights caused by out-of-distribution samples in the fine-tuning dataset. Then we ask the most important question- if we further increase the amplitude of those particular weight changes along the same direction, does it make the model more biased towards the out-of-distribution data? Later, in Section IV-C where we discuss our Maximizing Data Memorization (MDM) techniques, we illustrate that the answer to this question is ‘Yes’, and we can strategically focus on modifying the weights of these sensitive layers to enhance the model’s memorization of private data. But that is not all. With this also comes the salient question- If we amplify the weight changes in such a manner, does that affect the model utility? We also convincingly answer this question in Section IV-C by showing that such modification hardly impacts the model’s learning of regular English texts.
@@
We qualitatively define memorization $\mathbf{M}$ as a property of the language model in exposing particular training data points during inference. Quantitative measure of memorization $\mathbf{M}$ on $\mathcal{D_{\text{sen}}}$ includes but are not limited to data reconstruction [carlini2019secret], membership inference [mattern-etal-2023-membership], and exposure, which we discuss in detail in Section IV-F.
@@
TABLE I: A high-level taxonomy of our attack methods
IV Attack Methodology
IV-A Threat model
In our settings, both the adversary and the victim(s) are members of the client group participating in FL to fine-tune the LLM for a text auto-completion task, where they don’t share any gradient information with each other. We assume that the adversary has some context of the victim’s training data, specifically, the non-sensitive prefixes of the privacy-sensitive texts they want to extract from the victim’s local dataset. This assumption is highly practical in corporate and administrative contexts where communication relies on boilerplate templates. For example, automated notification emails (’Your verification code is…’), medical intake forms, or banking alerts follow rigid structural patterns. The adversary only needs to know the standard template (prefix) to extract the unique sensitive suffix. However, for the sake of a more inclusive threat model, we also perform our attacks with inaccurate prefixes, i.e., prefixes with perturbations. Apart from that, we consider two independent cases of the adversary’s capability: (1) the adversary is a stand-alone malicious client in the FL setting who does not collude with the server and observe the global model in each round, and (2) optionally, the malicious client can collaborate with the server which enables it to exploit a specific victim’s local model.